# CODES: Benchmarking Coupled ODE Surrogates

### v1

@@ Abstract
We introduce CODES, a benchmark for comprehensive evaluation of surrogate architectures for coupled ODE systems. Besides standard metrics like mean squared error (MSE) and inference time, CODES provides insights into surrogate behaviour across multiple dimensions like interpolation, extrapolation, sparse data, uncertainty quantification and gradient correlation. The benchmark emphasizes usability through features such as integrated parallel training, a web-based configuration generator, and pre-implemented baseline models and datasets. Extensive documentation ensures sustainability and provides the foundation for collaborative improvement. By offering a fair and multi-faceted comparison, CODES helps researchers select the most suitable surrogate for their specific dataset and application while deepening our understanding of surrogate learning behaviour.
@@ 1 Introduction and Overview
Coupled ordinary differential equations (ODEs) are ubiquitous in the natural sciences, and significant effort has gone into using machine learning models (surrogates) to replace the computationally expensive numerical methods required to solve such systems [4, 2, 5, 9, 3]. Despite individual benchmarking efforts when presenting such new surrogate architectures, there is a lack of comprehensive and fair comparisons across different architectures. This makes it difficult to identify the best surrogate for a specific task, especially since the optimal choice can depend on both the dataset and the target application.
The CODES benchmark (Coupled ODE Surrogates) addresses this gap by providing a framework for evaluating surrogates in the context of coupled ODE systems. Beyond accuracy, training, inference time, and computational cost, there are other crucial factors that differentiate architectures. As shown in section 3, surrogates vary in how they handle sparse data and perform in interpolation and extrapolation tasks. Some struggle more in highly dynamic regions, while uncertainty quantification is vital for hybrid approaches where numerical solvers may be needed for uncertain predictions. CODES highlights these differences, helping researchers and practitioners choose the best model for their task and deepening our understanding of whether surrogates truly capture the underlying dynamics of the data. The primary use cases include adding new datasets and selecting the most suitable surrogate or comparing new surrogates to baseline models.
CODES focuses on many essential aspects of benchmarking. Several baseline models and datasets are pre-implemented. Extensive documentation, featuring detailed doc-strings and type hints, ensures ease of use and project sustainability, supported by a dedicated documentation website [11]. To guarantee reproducibility, all models are trained and evaluated in a fully seeded manner. Usability is crucial for broad adoption, so features such as a web-based config-file generator, parallel training support, toggleable comparison modalities, and automated evaluation through tables and plots have been integrated. CODES is intended to grow as a collaborative project, with clear instructions on adding new datasets and surrogate models, encouraging community contributions.
@@ 2 Structure of the Benchmark
Datasets & Surrogates
CODES is largely inspired by astrochemistry, hence the two most complex datasets (osu2008 adapted from Grassi et al. [4] and branca24 from Branca & Pallottini [3]) come from this field. The remaining three datasets (lotka_volterra, simple_ode, and simple_reaction) are lower-dimensional baselines created by the authors of this paper. All datasets are hosted on Zenodo and automatically downloaded as needed. The current architectures include FullyConnected (FCNN, a standard fully-connected neural network), MultiONet (MON, an adaptation of DeepONet from [7] for multiple outputs as proposed in [8]), LatentNeuralODE (LNODE), and LatentPoly (LP, an autoencoder with a latent space neural ODE or a learnable latent space polynomial, both following [10]). All models receive as input the abundances at $t=0$ and then predict abundances for some later time $t>0$. Detailed explanations for all architectures and datasets are available in Appendix A and Appendix B as well as on the documentation website [11].
Modalities
@@
In addition to training a main model on the full dataset, the following modalities can be toggled to train additional models: Interpolation skips timesteps (or generally steps of the independent variable) in a specified interval. Extrapolation trains only on timesteps below a set threshold. Sparse thins out the training data by a specified factor. Batch trains the models with a user-defined batch size. By setting multiple values (e.g., intervals, thresholds), users can gain insights into and compare surrogate behaviour, answering questions such as: Which model handles sparse data best? Which performs the best interpolation? Uncertainty trains $N$ additional models with different random initializations and shuffled datasets to create a DeepEnsemble. For evaluation, users can configure additional output such as inference time, memory usage, loss curves, and different correlations. A configuration tool on the website [11] helps users generate the config.yaml.
Training & Benchmarking
@@
Training is the most time-intensive step in the benchmark since many models may need to be trained (e.g., 96 models in the evaluations presented in section 3). To accelerate this process, training is parallelized across all specified devices. Progress bars show the status of active jobs. Custom hyperparameters can be specified for each surrogate, and predefined optimal hyperparameters are available for the provided datasets. After training, evaluation results are generated as metrics and plots. These include individual (per-surrogate) and comparative results. Currently, up to 18 individual plots per surrogate and 12 comparative plots are provided. Metrics include standard measures like mean-squared error (MSE), mean absolute error (MAE), mean relative error (MRE), and inference time, as well as more advanced metrics, such as the Pearson correlation coefficient (PCC) between predictive uncertainty (via DeepEnsemble) and prediction error (PE), or between data gradients and PE. See Table 1 for a selection of the available metrics.
@@ 3 Results
The following results represent an exemplary evaluation of the benchmark. Models were trained on the osu2008 dataset, generated from a chemical reaction network with 224 reactions among 29 chemical species, simulated at constant temperature over 100 time steps. The dataset contains 1000 datapoints, split into 75% training, 5% validation, and 20% testing. Models were trained on NVIDIA TITAN Xp GPUs and evaluated on a NVIDIA GeForce RTX 2080 Ti. Optimal configurations for each model were determined using Optuna [1] and are detailed in Appendix A. While the results are specific to this dataset, broader patterns can be observed. Given the benchmark’s comprehensive nature, the following results are just a subset of the analyses and insights that can be drawn. Additional plots can be found in Appendix C or on the documentation website [11].
Performance
The main benchmark metrics, along with additional model parameters, are shown in Table 1. Despite significant differences between models, all show impressive performance — even the slowest model, LNODE, achieves speedups of at least three orders of magnitude compared to numerical methods (as discussed by Sulzer & Buck in [10]) and all models exhibit mean relative errors of below 4 %. MON performs best on accuracy metrics such as MSE, MAE and MRE, while LP has the fastest inference time, likely due to the computational simplicity of evaluating polynomials with fixed coefficients. Figure 1 displays the mean and median error of each surrogate over time, with spikes in the second half corresponding to more dynamic sections of the data.
@@
Interpolation, Extrapolation, and Sparsity
The differences in inductive bias between surrogates become more apparent under challenging conditions. Figure 2 shows the model MAE across three different benchmark modalities. In interpolation and extrapolation, FCNN outperforms all other models. The spikes observed in LNODE may result from intervals where the remaining timesteps are not representative of the full dataset, a challenge the other models seem to handle more effectively. In the sparse data setting, the autoencoder-type models, LNODE and LP, outperform the fully-connected models, FCNN and MON, suggesting greater robustness to reduced information.
@@
Uncertainty Quantification & Correlations
Comparing predictive uncertainty (the $1\sigma$ interval from DeepEnsemble predictions) with the mean absolute error over time in Figure 1, DeepEnsemble provides a generally reliable uncertainty estimate, particularly for LNODE and MON. This is also confirmed by comparing the MAE and mean uncertainty in Table 1. FCNN tends to be overconfident, underestimating the magnitude of its errors, while LP shows the opposite behavior. FCNN’s overconfidence is further reflected in its lowest PCC for UQ, computed across all corresponding uncertainty predictions and PE. Notably, the highest PCC between data gradients and PE occurs with LNODE, another indication that this architecture struggles more with non-smooth data than other surrogates.
@@ 4 Discussion
Limitations & Outlook.
Ensuring fair and representative comparisons can be challenging. CODES addresses this by seeding and training all models consistently within the benchmark. For the results in section 3, models were compared by training each until its loss trajectory stabilized, which is an effective but potentially imprecise approach. Alternatives such as training for a fixed number of epochs (implemented) or equal training time (not yet implemented) could also be argued for, and would significantly affect benchmark outcomes. Model performance heavily depends on hyperparameters, which were optimized via Optuna. This process is not yet integrated into the benchmark and must be performed manually; it may bias some models towards higher complexity, sacrificing speed. Architectural differences further complicate the comparison of the inference time. For example, LNODE is more efficient when predicting entire trajectories at once rather than generating individual predictions step-by-step, potentially skewing the comparison. We adopted the trajectory-based approach, but its representativeness depends on the target application. Finally, the benchmarking results are always dataset-specific. This implies that a broad range of datasets is needed to thoroughly assess surrogate performance. More importantly, findings are only reliable if they match the data distribution of the target application, emphasizing the need for representative training data.
CODES enables comprehensive comparisons of new and existing surrogate architectures and helps users select the ideal model for specific contexts, ensuring optimal performance in target applications. By improving surrogate selection, CODES contributes to faster and more reliable simulation frameworks, enabling higher resolution or more complex applications. Future updates will include integrated hyperparameter tuning, dataset-specific metrics and visualizations, and further exploration of inter- and extrapolation capabilities. A key focus will be adding new datasets and surrogates, for which we warmly invite community contributions, which will be credited appropriately.

### v2

@@ Abstract
We introduce CODES, a benchmark for comprehensive evaluation of surrogate architectures for coupled ODE systems. Besides standard metrics like mean squared error (MSE) and inference time, CODES provides insights into surrogate behaviour across multiple dimensions like interpolation, extrapolation, sparse data, uncertainty quantification, and gradient correlation. The benchmark emphasizes usability through features such as integrated parallel training, a web-based configuration generator, and pre-implemented baseline models and datasets. Extensive documentation ensures sustainability and provides the foundation for collaborative improvement. By offering a fair and multi-faceted comparison, CODES helps researchers select the most suitable surrogate for their specific dataset and application while deepening our understanding of surrogate learning behaviour.
@@ 1 Introduction and Overview
Coupled ordinary differential equations (ODEs) are ubiquitous in the natural sciences, and significant effort has gone into using machine learning models (surrogates) to replace the computationally expensive numerical methods required to solve such systems [4, 2, 5, 10, 3]. Despite individual benchmarking efforts when presenting such new approaches, there is a lack of comprehensive and fair comparisons across different architectures. This makes it difficult to identify the best surrogate for a specific task, especially since the optimal choice can depend on both the dataset and the target application.
The CODES benchmark (Coupled ODE Surrogates) addresses this gap by providing a framework for evaluating surrogates in the context of coupled ODE systems. Beyond accuracy, training, inference time, and computational cost, other crucial factors differentiate architectures. As shown in section 3, surrogates vary in how they handle sparse data and perform in interpolation and extrapolation tasks. Some struggle more in highly dynamic regions, while others may yield unreliable uncertainty estimates, which could be problematic in hybrid approaches where numerical solvers are employed for predictions with low certainty. CODES highlights these differences, helping researchers and practitioners choose the best model for their task and deepening our understanding of whether surrogates truly capture the underlying dynamics of the data. The primary use cases include adding a new dataset and identifying the most suitable surrogate for it or adding a new surrogate to the benchmark and comparing its performance to the baseline models on existing datasets.
CODES focuses on many essential aspects of benchmarking. Several baseline models and datasets are already implemented. Extensive documentation, featuring detailed doc-strings and type hints, ensures ease of use and project sustainability, supported by a dedicated documentation website [12]. To guarantee reproducibility, all models are trained and evaluated in a fully seeded manner. Usability is crucial for broad adoption. Therefore, features such as a web-based config-file generator, parallel training support, toggleable comparison modalities, and automated evaluation through tables and plots have been integrated. CODES is intended to grow as a collaborative project, with clear instructions on adding new datasets and surrogate models, encouraging community contributions.
@@ 2 Structure of the Benchmark
Datasets & Surrogates.
CODES is largely inspired by astrochemistry, hence the two most complex datasets come from this field: osu2008 was adapted from Grassi et al. [4] (an exemplary plot is shown in Figure 1) while branca24 was kindly provided by Branca & Pallottini from their recent paper [3]. The remaining three datasets (lotka_volterra, simple_ode, and simple_reaction) are lower-dimensional baselines created by the authors of this paper. All datasets are hosted on Zenodo and automatically downloaded as needed.
The following architectures are currently implemented:
  • FullyConnected (FCNN). A standard fully-connected neural network.
  • MultiONet (MON). An adaptation of DeepONet [8] for multiple outputs as proposed in [9].
  • LatentNeuralODE (LNODE). An autoencoder with a latent space neural ODE, following [11]
  • LatentPoly (LP). An autoencoder with a learnable latent space polynomial, also following [11].
All models receive as input the abundances at $t=0$ and then predict abundances for some later time $t>0$. Time is treated either as an additional input (FCNN, MON) or as a parameter that determines the length of the latent-space evolution (LNODE, LP). Detailed explanations for all architectures and datasets are available in Appendix A and Appendix B as well as on the documentation website [12].
Modalities.
In addition to training a main model on the full dataset, the following modalities can be toggled to train additional models:
  • Interpolation skips timesteps (or generally the independent variable) in a specified interval.
  • Extrapolation trains only on timesteps below a set threshold.
  • Sparse thins out the training data by a specified factor, using only every $n$-th sample.
  • Batch trains the models with a user-defined batch size.
  • Uncertainty trains $N$ additional models with different random initializations and shuffled datasets to create a DeepEnsemble [7] for uncertainty estimation.
@@
By setting multiple values (i.e., multiple intervals, thresholds, factors, or batch sizes), users can gain insights into and compare surrogate behaviour, answering questions such as: Which model handles sparse data best? Which performs the best interpolation? For evaluation, users can configure additional output such as inference time, memory usage, loss curves, and different correlations. A configuration tool on the website [12] helps users generate the config.yaml.
Training & Benchmarking.
@@
Training is the most time-intensive step in the benchmark, as for each setting an additional model is trained on the corresponding subset of training data (e.g., 96 models were trained for the evaluations presented in section 3). To accelerate this process, training is parallelized across all specified devices using a task-list approach. Progress bars show the status of active jobs. Custom hyperparameters can be specified for each surrogate, and predefined optimal hyperparameters are available for the provided datasets. After the training run finishes, evaluation results are generated in the form of metrics and plots. These include individual (per-surrogate) and comparative results. Currently, up to 18 individual plots per surrogate and 12 comparative plots are provided. Metrics include standard measures like mean-squared error (MSE), mean absolute error (MAE), mean relative error (MRE), and inference time. Additionally, there are more advanced metrics, such as the Pearson correlation coefficient (PCC) between predictive uncertainty (via DeepEnsemble) and prediction error (PE), or between data gradients and prediction error. See Table 1 for a selection of the available metrics.
@@ 3 Results
The following results represent an exemplary evaluation of the benchmark. Models were trained on the osu2008 dataset, generated from a chemical reaction network with 224 reactions among 29 chemical species, simulated at constant temperature over 100 time steps. The dataset contains 1000 data points, split into 75% training, 5% validation, and 20% testing. Models were trained on NVIDIA TITAN Xp GPUs and evaluated on an NVIDIA GeForce RTX 2080 Ti. Optimal configurations for each model were determined using Optuna [1] and are detailed in Appendix A. While the results are specific to this dataset, broader patterns can be observed. Given the benchmark’s comprehensive nature, the following results are just a subset of the analyses and insights that can be drawn. Additional plots can be found in Appendix C or on the documentation website [12].
Performance.
The main benchmark metrics, along with additional model parameters, are shown in Table 1. Despite significant differences between models, all show impressive performance — even the slowest model, LNODE, achieves speedups of at least three orders of magnitude compared to numerical methods (as discussed by Sulzer & Buck in [11]) and all models exhibit mean relative errors of below 4%. MON performs best on accuracy metrics such as MSE, MAE and MRE (followed closely by FCNN) , while LP has the fastest inference time, likely due to the computational simplicity of evaluating polynomials with fixed coefficients. Figure 2 displays the mean and median error of each surrogate over time. The spikes in the second half correspond to more dynamic sections of the data, with which some surrogates struggle more than others (cf. the section on UQ and correlations).
@@
Interpolation, Extrapolation, and Sparsity.
The differences in inductive bias between surrogates become more apparent under challenging conditions. Figure 3 shows the model MAE across three different benchmark modalities. In interpolation and extrapolation, FCNN outperforms all other models. The spikes observed in LNODE may result from intervals where the remaining timesteps are not representative of the full dataset, a challenge the other models seem to handle more effectively. In the sparse data setting, the autoencoder-type models, LNODE and LP, outperform the fully-connected models, FCNN and MON, for low numbers of training samples ($\leq 93$), suggesting greater robustness to settings with reduced information. This might be explained by the fact that the former involve a temporal evolution, which more closely mirrors the problem structure, while the latter treat time merely as an additional input.
@@
Uncertainty Quantification & Correlations.
Comparing the predictive uncertainty (the $1\sigma$ interval from DeepEnsemble predictions) with the mean absolute error over time in Figure 2 allows for insights into how well each surrogate architecture estimates its own prediction errors. DeepEnsemble seems to provide a generally reliable uncertainty estimate, particularly for LNODE and MON. This is also confirmed by the close coincidence of the MAE and mean uncertainty metrics in Table 1. FCNN tends to be overconfident, underestimating the magnitude of its errors, while LP slightly tends towards the opposite behavior. It could be argued that the magnitude of an uncertainty estimate is less important than its correlation to the errors, since the estimate could just be rescaled appropriately. Following this argument, we added the Pearson correlation coefficient (PCC) between the predicted uncertainty and the prediction error (calculated across all predictions in the test set) to the benchmark. This metric is lowest for FCNN, confirming the apparent overconfidence of this architecture. Additionally, we integrated the PCC between data gradients and prediction error. This metric is highest for LNODE, evidencing a proclivity towards higher errors at points with larger gradients. This provides an additional indication to our earlier impression that this architecture struggles more with non-smooth data than other architectures.
@@ 4 Discussion
Limitations.
Ensuring fair and representative comparisons can be challenging. CODES addresses this by seeding and training all models consistently within the benchmark. The limitations of the benchmark can roughly be divided into two categories: the influence of architectural differences on comparability and the data-dependency of the results.
For the results in section 3, models were compared by training each until its loss trajectory stabilized, which is an effective but potentially imprecise approach. Alternatives such as training for a fixed number of epochs (implemented) or equal training time (not yet implemented) could also be argued for, and would significantly affect benchmark outcomes. Model performance heavily depends on hyperparameters, which were optimized via Optuna. This process is not yet integrated into the benchmark and must be performed manually; it may bias some models towards higher complexity, sacrificing speed. Architectural differences further complicate the comparison of the inference time. For example, LNODE is more efficient when predicting entire trajectories at once rather than generating individual predictions step-by-step, potentially skewing the comparison. We adopted the trajectory-based approach, but its representativeness depends on the target application.
In principle, training surrogates to model an ODE is an infinite-data problem because the ground truth data is generated with a numerical solver from randomly sampled initial conditions. In addition to allowing for the generation of arbitrary amounts of training data, this problem structure also invites online or active learning approaches. However, the benchmark currently only supports finite datasets provided in hdf5 format. Finally, the benchmarking results are always dataset-specific. This implies that a broad range of datasets is needed to thoroughly assess surrogate performance and necessitates further insights into each dataset to inform comparisons between them. Additionally, findings are only reliable if the distribution of the training data matches that of the target application, emphasizing the need for representative training data.
Outlook.
CODES enables comprehensive comparisons of new and existing surrogate architectures and helps users select the ideal model for their specific context, ensuring optimal performance in target applications. By improving surrogate selection, CODES contributes to faster and more reliable simulation frameworks, enabling higher resolution or more complex applications. Future updates will include integrated hyperparameter tuning, dataset-specific metrics and visualizations, further exploration of inter- and extrapolation capabilities, additional baseline models, and potentially support for online training. A key focus will be adding new datasets and surrogates, for which we warmly invite community contributions that will be credited appropriately.
Acknowledgments and Disclosure of Funding
This work is funded by the Carl-Zeiss-Stiftung through the NEXUS program and supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1 - 390900948 (the Heidelberg STRUCTURES Cluster of Excellence) via the exploratory project EP10.2.